### v1

# Length independent generalization bounds for deep SSM architectures with stability constraints

Abstract
Many state-of-the-art models trained on long-range sequences, for example S4, S5 or LRU, are made of sequential blocks combining State-Space Models (SSMs) with neural networks. In this paper we provide a PAC bound that holds for these kind of architectures with stable SSM blocks and does not depend on the length of the input sequence. Imposing stability of the SSM blocks is a standard practice in the literature, and it is known to help performance. Our results provide a theoretical justification for the use of stable SSM blocks as the proposed PAC bound decreases as the degree of stability of the SSM blocks increases.
Keywords: generalization bounds, state-space models, stability, Rademacher complexity
1 Introduction
The demand for learning powerful representation of time series have been present for decades. The problem of modeling long-range sequences, i.e. sequences with large number of time-steps, is an especially challenging task of the field. Recurrent Neural Networks (RNNs) have always had a hard time dealing with long time series, even advanced architectures such as LSTM or GRU networks did not make a breakthrough. Transformer based architectures managed to imrpove the results in some cases, however still performed poorly on more difficult tasks.
Recently, several novel architectures have been published that are outperforming previous models by a significant margin. The common basis of these models are the so-called Structured State-Space Models (SSMs), which are basically dynamical systems of either continuous or discrete time. Some of the well-performing architectures are the S4 and S5 models and their variants, or another linear system based model called LRU. These models all contain some form of SSMs, either linear or nonlinear, and combine them with some sort of nonlinear transformation from a single nonlinearity to deep neural networks. These blocks of SSMs and neural networks then may be concatenated resulting in a deep structure.
These trending architectures are usually validated empirically on commonly accepted datasets such as the Long Range Arena benchmark. On the other hand, the theoretical foundations of deep, SSM based structured are yet to be discovered. One key point of these models is that they are equipped with some form of stability constraints. For example, the LRU model contains linear SSMs with a certain exponential parametrization of its coefficient matrices enforcing stable behavior of the system. This motivates the question:
What is the role of stability in the success of deep SSM architectures for long-range sequences?
In this paper, we focus on this question and provide a theoretical framework to analyze the model’s generalization behavior in a rigorous manner. We show that stability of deep SSM architectures has an influence on their Rademacher complexity, resulting in a generalization bound that does not depend on the length of the input sequence. In more details:
  • We prove that the Rademacher complexity of a set of SSMs can be upper bounded by the maximal system norm of the considered systems. We consider two types of system norms: $H_{2}$ and $\ell_{1}$ norm. These norms defined for dynamical systems are well-known in control theory, along with their connection to quadratic stability. This result clearly suggests that enforcing stability at least up to a certain level is far beyond being a practical necessity and, in fact, a significant aspect of understanding the essence of these architectures. Consequently, we consider this point as the main takeaway of our paper.
  • Furthermore, we establish an upper bound on the Rademacher complexity of multilayer models by introducing the property of Rademacher Contraction for various components of deep SSM structures, containing multiple separate SSMs combined with nonlinear transformations in between, which is applicable for the popular architectures mentioned above.
  • Using the previous results we establish a PAC-bound on the generalization error of deep SSMs. The resulting bound does not depend on the length of the input sequences, and it only implicitly depends on the depth of the model.
Most of the state-of-the art deep SSM architectures for long range sequences are covered by our results. Moreover, all our our definitions, theorems and proofs hold for both continuous and discrete time systems, as well as for both classification and regression tasks. For the sake of compactness, we stick to the discrete time representations.
Outline of the paper. In Section 2 we present the related literature, then we set notations, definitions and our assumptions related to deep SSMs in Section 3. We propose our main result in Section 4. The majority of the proofs and some additional details are shown in the Appendix.
2 Related work.
There are several PAC bounds for either discrete or continuous-time RNNs in Koiran and Sontag (1998); Sontag (1998); Hanson et al. (2021) by using VC dimension usually through covering numbers. PAC bounds for RNNs based on Rademacher complexity were presented in Wei and Ma (2019); Akpinar et al. (2020); Joukovsky et al. (2021); Chen et al. (2020), while in Zhang et al. (2018) the authors developed PAC-Bayesian bounds. Bounds for general RNNs are related to SSMs as Linear Time-Invariant (LTI) dynamical systems are essential elements almost for all SSMs and they are a special class of RNNs. However all of the previously mentioned PAC bounds depend on the number of time steps and the error bounds tend to infinity with the integration time in continuous case or with the number of time steps in discrete case. That is, these error bounds are not meaningful for prediction errors on long-range sequences.
PAC bounds for both linear and nonlinear dynamical systems are either consider problems not applicable to LTI systems (therefore not applicable for SSMs considered in this paper, see Alquier and Wintenberger (2012); Alquier et al. (2013); Shalaeva et al. (2020); Haussmann et al. (2021); Marion (2023); Haddouche and Guedj (2022)), or not length independent (Hanson and Raginsky (2024)). For example PAC bounds were developed for discrete-time autoregressive models in Alquier and Wintenberger (2012); Alquier et al. (2013); Shalaeva et al. (2020), for stochastic differential equations in Haussmann et al. (2021), for neural ordinary differential equations in Marion (2023), or for online learning in Haddouche and Guedj (2022). In Hanson and Raginsky (2024) the authors propose a PAC bound based on Rademacher complexity for input-affine non-linear systems however their bound is still exponential in the length of the integration interval.
The generalization bound for single vanilla RNNs in (Chen et al., 2020, Theorem 2) is an upper bound of the $H_{1}$ norm proposed in this paper, see e.g. Chellaboina et al. (1999), thus our results based on the $H_{2}$ norm is tighter. Their bound is consistent with our in case of a single linear RNN block. Additionally, they assume bounded matrix norms and their bound depends on the dimension of the representation. In comparison our bound does not depend on the dimension of the state space and we assume only that the systems have bounded $H_{2}$ norm, even in case of stacked RNNs.
In Trauger and Tewari (2024) the authors propose a sequence length independent Rademacher complexity bound for a single layer transformer architecture. For multi layer transformers they improve slightly the result in Edelman et al. (2022) however the bound grows logarithmically with the sequence length.
There are some relevant results outside of the field of RNNs or dynamical systems dealing with the depth of deep neural networks in generalization bounds Bartlett et al. (2017); Liang et al. (2019); Golowich et al. (2018); Truong (2022b). In Golowich et al. (2018) the authors derived a depth independent bound under the condition of bounded Schatten p-norm and a bound with polynomial dependence on depth for Rademacher complexity for DNNs with ReLU activations by applying contraction. Similar approaches utilized spectral norm Bartlett et al. (2017) or Fisher-Rao norm Liang et al. (2019) to establish bounds with non exponential dependence on depth. In a recent paper Truong (2022b) the author extends the bound in Golowich et al. (2018) for non ReLU activations and show that the new, non-vacuous bound is depth independent.
3 Problem setup
In this section, we formally define deep SSM models in way that popular architectures from the literature are contained, the generalization error and the concept of PAC bounds. As our results stand for both continuous-time (CT) and discrete-time (DT) SSMs, in some cases we introduce parallel definitions corresponding to each case.
Notation. We denote scalars with lowercase characters, vectors with lowercase bold characters and matrices with uppercase characters. For a matrix $A$ let $A_{i}$ denote its $i$th row. The symbol $\odot$ denotes the elementwise product. We use $[n]$ to denote the set $\{1,2,\ldots,n\}$ for $n\in\mathbb{N}$.
For vector valued time dependent functions originating from a finite set, the notation $\mathbf{u}_{i}^{(j)}(t)$ refers to the $j$th coordinate of the $i$th function at time $t$. Similarly, in the discrete time case, the notation $\mathbf{u}_{i}^{(j)}[k]$ refers to the $j$th component of the $i$th time series at time step $k$.
$\Sigma$ denotes a dynamical system specified in the context. The constant $n_{\text{in}}$ refers to the dimension of the input sequence, $T$ refers to its length in time, while $n_{\text{out}}$ is the dimension of the output (not necessarily a sequence).
Denote by $\ell^{2,2}_{T}(\mathbb{R}^{n})$ and $\ell^{\infty,\infty}_{T}(\mathbb{R}^{n})$ the Banach spaces generated by the all finite sequences over $\mathbb{R}^{n}$ of length $T$ with the norm $\left\lVert\mathbf{u}\right\rVert^{2}_{\ell^{2,2}_{T}(\mathbb{R}^{n})}=\sum_{k=0}^{T-1}\left\lVert\mathbf{u}[k]\right\rVert_{2}^{2}$ and $\left\lVert\mathbf{u}\right\rVert_{\ell^{\infty,\infty}_{T}(\mathbb{R}^{n})}=\sup\limits_{k=0,\ldots,T-1}\left\lVert\mathbf{u}[k]\right\rVert_{\infty}$ respectively. More generally, denote by $\ell^{2,2}(\mathbb{R}^{n})$ and $\ell^{\infty,\infty}(\mathbb{R}^{n})$ the Banach spaces generated by the all inifinite sequences over $\mathbb{R}^{n}$ such that the quantities $\left\lVert\mathbf{u}\right\rVert^{2}_{\ell^{2,2}(\mathbb{R}^{n})}=\sum_{k=0}^{\infty}\left\lVert\mathbf{u}[k]\right\rVert_{2}^{2}$ and $\left\lVert\mathbf{u}\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n})}=\sup\limits_{k}\left\lVert\mathbf{u}[k]\right\rVert_{\infty}$ are well defined and finite. If $\mathbf{u}=\mathbf{u}[0]\ldots,\mathbf{u}[T-1]$ is a finite sequence of length $T$, then we can interpret it as an infinite sequence $\mathbf{u}=\mathbf{u}[0]\ldots,\mathbf{u}[T-1],0,0,\ldots$; elements of which are zero after the $T$th element. For a Banach space $\mathcal{X}$, $B_{\mathcal{X}}(r)=\{x\in\mathcal{X}\mid\left\lVert x\right\rVert_{\mathcal{X}}\leq r\}$ denotes the ball of radius $r>0$ centered in zero.
We use the symbols $E_{(u,y)\sim\mathcal{D}}$, $P_{(u,y)\sim\mathcal{D}}$, $E_{\mathbf{S}\sim\mathcal{D}^{N}}$ and $P_{\mathbf{S}\sim\mathcal{D}^{N}}$ to denote expectations and probabilities w.r.t. a probability measure $\mathcal{D}$ and its $N$-fold product $\mathcal{D}^{N}$ respectively (see Section 3.2). The notation $\mathbf{S}\sim\mathcal{D}^{N}$ tacitly assumes that $\mathbf{S}\in(\mathcal{U}\times\mathcal{Y})^{N}$, i.e. $\mathbf{S}$ is made of $N$ triplets of input and output trajectories.
3.1 Deep SSMs
A State-Space Model (SSM) is a discrete-time linear dynamical system of the form
$\Sigma\begin{cases}&\mathbf{x}[k]=A\mathbf{x}[k-1]+B\mathbf{u}[k],\leavevmode\nobreak\ \mathbf{x}[0]=0\\ &\mathbf{y}[k]=C\mathbf{x}[k]+D\mathbf{u}[k]\end{cases}$ (1)
where $A\in\mathbb{R}^{n_{x}\times n_{x}},B\in\mathbb{R}^{n_{x}\times n_{u}},C\in\mathbb{R}^{n_{y}\times n_{x}}$ and $D\in\mathbb{R}^{n_{y}\times n_{u}}$ are matrices, is some function of $\mathbf{u}[k],\mathbf{x}[k]$, and $k=1,2,\ldots,T$, where $T$ is the number of time steps. We consider the value of $T$ to be fixed. The reason behind this is rather technical, namely to conviniently handle the Pooling layer (see Definition 6). We emphasize that the generalization bound in Theorem 18 is independent of $T$. For regression tasks the model does not contain any pooling layer, thus this restriction is not needed.
Input-output maps of SSMs. The SSM (1) can be run for finite or infinite number of time steps, and hence it realizes a sequence to sequence transformation. SSM induce as sequence to sequence transformations: they map any input sequence $\mathbf{u}$ of any length (including infinite length) to the unique output sequence $\mathbf{y}$ of the same length, i.e. an SSM $\Sigma$ induces the input-output map of $\Sigma$, denoted by $\mathcal{S}_{\Sigma}$. Moreover, $\mathcal{S}_{\Sigma}$ is causal, $\mathbf{y}[t]$ depends only on the first $t+1$ inputs. The outputs of SSMs are linear combinations of past inputs, in fact, input-output maps can be expressed by a convolution $\mathbf{y}[k]=\mathcal{S}_{\Sigma}(\mathbf{u})[k]=\sum_{j=0}^{k}H_{j}\mathbf{u}(k-j)$, where $H_{0}=D$ and $H_{j}=CA^{j-1}B$, $j>0$. That is, SSMs mix inputs along the time axis, while preserving causality
Stability and input-output maps of SSMs as operators on $\ell^{p,p}$, $p=\infty,2$. For SSMs to be robust transformations of sequences to sequences, one needs stability. Stability is one, if not the most important concept used in dynamical systems and control theory Antoulas (2005). Intuitively, the solutions of a stable system are continuous in the initial state and input. In particular, for stable systems, a small perturbation in past inputs will not result in an increasing error in future outputs.
Definition 1 (Antoulas (2005))
SSM of the form (1) is internally stable, if the matrix $A$ is Schur, meaning all the eigenvalues of $A$ are inside the complex unit disk.
In particular, a sufficient (but not necessary) condition for stability is that $A$ is a contraction, i.e. $\|A\|_{2}<1$.
A stable SSM $\Sigma$ is not only robust to perturbations, but it is well-known Antoulas (2005) that its input-output map $\mathcal{S}_{\Sigma}$ acts as a linear operator $\ell^{p,p}(\mathbb{R}^{n_{u}})\to\ell^{r,r}(\mathbb{R}^{n_{y}})$, for any choice $p,r\in\{\infty,2\}$, as for any $\mathbf{u}\in\ell^{p,p}(\mathbb{R}^{n_{u}})$, $\mathcal{S}_{\Sigma}(\mathbf{u})\in\ell^{r,r}(\mathbb{R}^{n_{y}})$. In particular, $\mathcal{S}_{\Sigma}$ has a well-defined induced norm as a linear operator, defined in the usual way, $\left\lVert\Sigma\right\rVert_{p,r}=\sup_{\mathbf{u}\in\ell^{p,p}(\mathbb{R}^{n_{u}})}\frac{\left\lVert\mathcal{S}_{\Sigma}(\mathbf{u})\right\rVert_{\ell^{r,r}(\mathbb{R}^{n_{y}})}}{\left\lVert\mathbf{u})\right\rVert_{\ell^{r,r}(\mathbb{R}^{n_{y}})}}$. In this paper, we will use the induced norms $\left\lVert\Sigma\right\rVert_{2,\infty}$ and $\left\lVert\Sigma\right\rVert_{\infty,\infty}$, which can be upper bounded by the following two norms defined on SSMs.
Definition 2 (Chellaboina et al. (1999))
For a SSM $\Sigma$ of the form (1) define the $\ell_{1}$ and $H_{2}$ norm of $\Sigma$, denoted by $\left\lVert\Sigma\right\rVert_{1}$ and $\left\lVert\Sigma\right\rVert_{2}$ respectively,
$\left\lVert\Sigma\right\rVert_{1}:=\max\limits_{1\leq i\leq n_{y}}\left\lVert D_{i}\right\rVert_{1}+\sum\limits_{k=0}^{\infty}\left\lVert C_{i}A^{k}B\right\rVert_{1},\quad\left\lVert\Sigma\right\rVert_{2}:=\sqrt{\left\lVert D\right\rVert_{F}^{2}+\sum\limits_{k=0}^{\infty}\left\lVert CA^{k}B\right\rVert_{F}^{2}}$
Lemma 3 (Chellaboina et al. (1999))
For a system of form (1) $\left\lVert\Sigma\right\rVert_{\infty}\leq\left\lVert\Sigma\right\rVert_{1}$ and $\left\lVert\Sigma\right\rVert_{2,\infty}\leq\left\lVert\Sigma\right\rVert_{2}$.
The norms defined above will play a cruicial role in in the main result of the paper, as they will allow us to bound the Rademacher complexity of the deep SSM model.
Remark 4 (Computing $\left\lVert\Sigma\right\rVert_{i}$,$i=1,2$)
The norm $\left\lVert\Sigma\right\rVert_{2}$ can be computed by solving Sylvester equations, for which standard numerical algorithms exist Antoulas (2005). The computation of $\left\lVert\Sigma\right\rVert_{1}$ is more involved, it can be computed by taking a sufficiently large finite sum instead of the infinite sum used in its definition. If $\|A\|_{2}<\beta<1$, then an easy calculation reveals that $\left\lVert\Sigma\right\rVert_{1}\leq\|D\|_{2}+\frac{\|B\|_{2}\|C\|_{2}}{1-\beta}$ and $\left\lVert\Sigma\right\rVert_{2}\leq\sqrt{\|D\|_{F}^{2}+\frac{n_{y}\|B\|_{2}^{2}\|C\|_{2}^{2}}{1-\beta^{2}}}$.
Relationship with continuous-time models. In the literature, SSMs are often derived by discretizing a continuous-time linear differential equation in time (e.g. Gu and Dao (2023) and references therein). If the discretization step $\Delta$ is a fixed constant, then we obtain time-invariant linear system of the form (1). In some models, the discretization step can depend on the input and state of the continuous-time state-space representation. In the latter case, one still obtains a linear discrete-time state-space representation, but then the matrices describing the system equation depend on the current input and state, and hence the matrices are time-dependent. In this paper we will stick to time-invariant models (1), for the sake of simplicity.
Deep SSM models. In this paper, we focus our attention to deep SSM models comoposed of multiple linear SSMs. The main reason is that most of the state-of-the-art architectures for long-range sequences are based on linear system. A key characteristics of deep SSM models are the combination of SSM layers with nonlinear transformations, typically some kind of neural network that is constant in time. The general definition we use is the following.
Definition 5
A DT-SSM block (or simply SSM block) is a function $f^{\text{DTB}}:\ell^{p,q}(\mathbb{R}^{n_{u}})\to\ell^{r,s}(\mathbb{R}^{n_{u}})$ that is composed of a stable SSM followed by a nonlinear transformation that is constant in time. That is, $f^{\text{DTB}}(\mathbf{u})[k]=g(\mathcal{S}_{\Sigma}(\mathbf{u})[k])+\alpha\mathbf{u}[k]$ for some $g:\mathbb{R}^{n_{u}}\to\mathbb{R}^{n_{u}}$ for all $k\in\left[T\right]$. The function $g$ is represented as either an MLP or a GLU network (see Definitions 7 and 8).
We incorporate $\alpha$ so that the definition covers residual connections (typically $\alpha$ is either 1 or 0). The previous definition is inspired by the series of popular architectures mentioned in the introduction. Similarly, the definition of deep SSM models is also based on these architectures. A deep SSM model consists of SSM blocks along with an encoder, and a decoder transformation preceded by a time-pooling layer. In practice, it is common to apply some normalization techniques, such as batch or layer normalization. As they are not essential for our result, we omit them from the definition. Indeed, once training is finished, a normalization layer corresponds to applying a neural network with linear activation function, i.e., it can be integrated into one of the neural network layer. Since the objective of PAC bounds is to bound the generalization error for already trained models, for the purposes of PAC bounds normalization layers can be viewed as an additional neural network layer.
Definition 6
A discrete time deep SSM model for classification is a function $f:\ell^{p,q}(\mathbb{R}^{n_{\text{in}}})\to\mathbb{R}^{n_{\text{out}}}$ of the form $f=f^{\text{Dec}}\circ f^{\text{Pool}}\circ f^{\text{B}_{L}}\circ\ldots\circ f^{\text{B}_{1}}\circ f^{\text{Enc}}$. The functions $f^{\text{Enc}}$ and $f^{\text{Dec}}$ are linear transformations which are constant in time, while $f^{\text{B}_{i}}$ is a DT-SSM block for all $i$. By pooling we mean the operation $f^{\text{Pool}}(\mathbf{u})=\frac{1}{T}\sum_{k=1}^{T}\mathbf{u}[k]$, an avarage pooling over the time axis.
For regression tasks, we omit the pooling layer and define the Decoder as a linear transformation that is identical for all time steps.
Model SSM Block
S4 Gu et al. (2021) LTI, $A=\Lambda-PQ^{*}$ block-diagonal SSM + nonlinear activation
S4D Gu et al. (2022) LTI, $A=-\text{exp}(A_{Re})+i\cdot A_{Im}$ block-diagonal SSM + nonlinear activation
S5 Smith et al. (2022) LTI, diagonal $A$ SSM + nonlinear activation
LRU Orvieto et al. (2023) LTI, diagonal $A$ complex exponential parametrization SSM + MLP skip connection
Table 1: Summary of popular deep SSM models.
The Encoder and Decoder layers are given by the weight matrices $W^{\text{Enc}}$ and $W^{\text{Dec}}$. Therefore $f^{\text{Enc}}(\mathbf{u})[k]=W^{\text{Enc}}\mathbf{u}[k]$ and $f^{\text{Dec}}(\mathbf{u})[k]=W^{\text{Dec}}\mathbf{u}[k]$ for all $k\in[T]$. We use the slightly abused notations $f^{\text{Enc}}\equiv\langle W^{\text{Enc}},\cdot\rangle$ and $f^{\text{Dec}}\equiv\langle W^{\text{Dec}}\cdot\rangle$.
As for the Neural Network components of an SSM block, we consider the following two variants.
Definition 7 (MLP layer)
An MLP layer is a function from $\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})$ to $\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})$ that is induced by applying a deep neural network $f:\mathbb{R}^{n_{u}}\to\mathbb{R}^{n_{u}}$ for each timestep. A neural network of $L$ layer is a function of the form $f=f_{W_{1},\mathbf{b}_{1}}\circ\ldots\circ f_{W_{L},\mathbf{b}_{L}}\circ g_{W_{L+1},\mathbf{b}_{L+1}}$, where $f_{W,\mathbf{b}}(\mathbf{x})=\rho(g_{W,\mathbf{b}}(\mathbf{x}))$ is called a hidden layer, $g_{W,\mathbf{b}}(\mathbf{x})=W\mathbf{x}+\mathbf{b}$ is called preactivation and $\rho$ is the activation function, which is identical for all layers of the network and is either sigmoid or ReLU. The matrices are of the size $W_{i}\in\mathbb{R}^{n_{i+1}\times n_{i}}$ and $\mathbf{b}\in\mathbb{R}^{n_{i}}$ such that $n_{1}=n_{u}$ and $n_{L+1}=n_{u}$. By slight abuse of notation, for $\mathbf{u}\in\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})$ let $f(\mathbf{u})\in\ell^{\infty,\infty}(\mathbb{R}^{n_{v}})$ such that $f(\mathbf{u})[k]=f(\mathbf{u}[k])$ for all $1\leq k\leq T$.
Definition 8 (GLU layer Smith et al. (2022))
A GLU layer is a function of the form $GLU:\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})\to\ell^{\infty,\infty}(\mathbb{R}^{n_{v}})$ parametrized by a linear operaetor $W$ such that $GLU(\mathbf{u})[k]=GELU(\mathbf{u}[k])\odot\sigma(W(GELU(\mathbf{u}[k])))$, where $\sigma$ is the sigmoid function and GELU is the Gaussian Error Linear Unit Hendrycks and Gimpel (2016).
Note, that this definition of GLU layer differs from the original definition in Dauphin et al. (2017), because in deep SSM models GLU is usually applied individually for each time step, without any time-mixing operations. See Appendix G.1 in Smith et al. (2022). The linear operation $W$ is usually represented by a convolution operation.
3.2 Learning problem
We consider the usual supervised learning framework for sequential input data. The considered models, parametrized by $\theta$, are of the form $f_{\theta}:\mathcal{U}^{+}\to\mathcal{Y}$, where $\mathcal{U}\subseteq\mathbb{R}^{n_{\text{in}}}$ is the input space, $\mathcal{U}^{+}$ denotes the set of all finite sequences and $\mathcal{Y}$ is the output space, either a finite set for classification or $\mathcal{Y}\subseteq\mathbb{R}^{n_{\text{out}}}$ for regression. Without loss of generality we assume $n_{\text{out}}=1$. In practice, $\theta$ is usually obtained by some learning algorithm, such as Gradient Flow or (Stochastic) Gradient Descent. In this paper, we are agnostic regarding the origin of $\theta$ as we prove a generalization bound that holds for a set of models, therefore we omit the subscript $\theta$ from the notation of $f$.
For training we will use input sequence of the same length $T$. More precisely, let us denote by $\mathcal{U}^{T}$ the set of all sequences of elements of $\mathcal{U}$ of length $T$. A dataset is an i.i.d sample of the form $S=\{(\mathbf{u}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$ where $\mathbf{u}_{i}\in\mathcal{U}^{T}$ and $\mathbf{y}_{i}\in\mathcal{Y}$ for all $i$, where the random sample is with respect to some probability measure $\mathcal{D}$ defined on the $\sigma$-algebra generated by the Borel sets of $\mathcal{U}^{T}\times\mathcal{Y}$, where $\mathcal{U}^{T}\times\mathcal{Y}$ is considered with respect to the standard topology as a subset of $(\mathbb{R}^{n_{\text{in}}})^{T}\times\mathbb{R}^{n_{\text{out}}}$.
An elementwise loss function is of the form $\ell:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ and is assumed to be $K_{\ell}$-Lipschitz-continuous, i.e. $|\ell(y_{1},y_{1}^{\prime})-\ell(y_{2},y_{2}^{\prime})|\leq K_{\ell}(|y_{1}-y_{2}|+|y_{1}^{\prime}-y_{2}^{\prime}|)$ for all $y_{1},y_{2},y_{1}^{\prime},y_{2}^{\prime}\in\mathbb{R}$, as well as $\ell(y,y)=0$ for all $y\in\mathbb{R}$. Let $\mathcal{L}_{emp}^{S}(f)=\frac{1}{N}\sum_{i=1}^{N}l(f(\mathbf{u}_{i}),y_{i})$ denote the empirical loss of a model $f$ w.r.t a dataset $S$. We denote the true error by $\mathcal{L}(f)=\mathbf{E}_{(\mathbf{u},y)\sim\mathcal{D}}[\ell(f(\mathbf{u}),y)]$. The generalization error or gap of a model $f$ is defined as $|\mathcal{L}_{emp}^{S}(f)-\mathcal{L}(f)|$. In practice, we can approximate the generalization gap by the empirical generalization gap, i.e. the loss difference on the training data and some test data (see Devroye et al. (2013)).
3.3 Assumptions
Before moving forward to discuss the main result, we summarize the assumptions we make in the paper for the sake of readibility. Hereinafter we denote by $\mathcal{F}$ a set of deep SSM models represented by its direct product of its layerwise parameters. Furthermore, let $\mathcal{E}$ denote the set of all SSM models $\Sigma$ for which there is a model $f\in\mathcal{F}$ such that $\Sigma$ is an SSM layer of $f$.
Assumption 9
We assume the following properties hold.
  1. 1.
    Scalar output. Let $n_{\text{out}}=1$.
  2. 2.
    Lipschitz loss function. Let the elementwise loss $l$ be $L_{l}$-Lipschitz continuous.
  3. 3.
    Bounded input. There exist $K_{\mathbf{u}}>0$ and $K_{y}>0$ such that for all input trajectory $\mathbf{u}$ we have $\left\lVert\mathbf{u}\right\rVert_{\ell^{2,2}(\mathbf{R}^{n_{\text{in}}})}\leq K_{\mathbf{u}}$ and for all labels we have $|y|\leq K_{y}$
  4. 4.
    Stability. All $\Sigma\in\mathcal{E}$ are internally stable, implying $\left\lVert\Sigma\right\rVert_{p}<+\infty$ for $p=1,2$. Therefore we assume there exist $K_{p}>0$ such that $\sup\limits_{\Sigma\in\mathcal{E}}\left\lVert\Sigma\right\rVert_{p}<K_{p}$ for $p=1,2$.
  5. 5.
    Bounded Encoder and Decoder. We assume the Encoder and Decoder have bounded operator norms, i.e. $\sup\limits_{W^{\text{Enc}}\in\mathcal{W}_{\text{Enc}}}\left\lVert W^{\text{Enc}}\right\rVert_{2,2}<K_{\text{Enc}}$ and $\sup\limits_{W^{\text{Dec}}\in\mathcal{W}_{\text{Dec}}}\left\lVert W^{\text{Dec}}\right\rVert_{\infty,\beta}<K_{\text{Dec}}$ for $\beta\in\mathbb{N}\cup\{\infty\}$.
Assumption 1. is not restrictive as we consider classification.
The Lipschitzness in Assumption 2. holds for most of the loss functions used in practice. We mention that even the square-loss is Lipschitz on a bounded domain. From the practical aspect, the upped boundedness is also mild, as parameters along the learning algorithm’s trajectory usually make $l$ bounded. In the worst case, $l$ is bounded on a bounded domain due to being Lipschitz.
Assumption 3. is yet again standard in the literature. Even in practical applications the input is usually normalized or standardized as a preprocessing step before learning.
Assumption 4. is the most important one as it plays a central role in our work. The motivation behind this assumption is twofold. First, in practical implementation of SSM based architectures, it is very common to apply some structured parametrization of the matrices of the systems, which leads to learning stable matrices. In many cases, the underlying intention is numerical stability of the learning algorithm, however we argue that the major advantage of such parametrizations is to ensure a stable behavior of the system. Second, similar stability assumptions are standard in control theory.
Assumption 5. is again fairly standard, as it requires that the weights of the encoder and decoder are bounded.
4 Main results
Our main result is a Rademacher complexity based generalization bound for deep SSM models that does not depend on the sequence length. The main challenge to establish such bounds are threefold:
  1. 1.
    without any assumption on the SSMs, their Rademacher complexity itself is not trivial to upper bound,
  2. 2.
    even if one had a bound on the Rademacher complexity of SSMs, it needs to be extened to a block of SSMs and deep neural networks,
  3. 3.
    deep SSM structures usually contain several of these blocks, therefore any estimation on the Rademacher complexity of the model class has to address this situation.
As to the first point, we argue that the stability of the SSMs has a crucial role in the generalization ability of these kind of models. While applying stable parametrization is common in practical implementations, to best of our knowledge, the literature has been lacking any theoretical guarantees on the effect of stability to the model’s performance. To this end, we show that the Rademacher complexity of a set of SSMs can be upper bounded by a term which has a tight dependence on the maximal $H_{2}$ norm of the SSMs.
The second and third points can be translated to the task of estimating the Rademacher complexity of a deep structure, where each block of the composition may have a different mathematical nature. We can overcome this obstacle by introducing a property of functions, referred to as Rademacher Contraction, that is universal enough to include functions represented by both SSMs and neural networks.
Definition 10 ($(\mu,c)$-Rademacher Contraction)
Let $X_{1}$ and $X_{2}$ be subsets of Banach spaces $\mathcal{X}_{1},\mathcal{X}_{2}$, with norms $\|\cdot\|_{\mathcal{X}_{1}}$ and $\|\cdot\|_{\mathcal{X}_{2}}$, and let $\mu\geq 0$ and $c\geq 0$. A set of functions $\Phi=\{\varphi:X_{1}\to X_{2}\}$ is said to be $(\mu,c)$-Rademacher Contraction, or $(\mu,c)$-RC in short, w.r.t. $\left\lVert\cdot\right\rVert_{X_{2}}$ and $\left\lVert\cdot\right\rVert_{X_{1}}$, if for all $n\in\mathbb{N}^{+}$ and $Z\subseteq X_{1}^{n}$ we have
$\mathbb{E}_{\sigma}\left[\sup\limits_{\varphi\in\Phi}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{n}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\varphi(\mathbf{u}_{i})\right\rVert_{\mathcal{X}_{2}}\right]\leq\mu\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{n}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\right\rVert_{\mathcal{X}_{1}}\right]+\frac{c}{\sqrt{N}},$ (2)
where $\sigma_{i}$ are i.i.d. Rademacher random variables, $i\in[N]$, i.e. $\mathbb{P}(\sigma_{i}=1)=\mathbb{P}(\sigma_{i}=-1)=1/2$.
Rademacher Contractions in the literature. Special cases of the inequality in Definition 10 can be found in the literature. In Golowich et al. (2018) the authors considered biasless ReLU networks and proved a similar inequality using Talagrand’s contraction lemma Ledoux and Talagrand (1991). In Truong (2022b), the author considered neural networks with dense and convolutional layers and derived a PAC bound via bounding the Rademacher complexity. One of the key technical achievements is Theorem 9, which is a more general version of the inequality in Golowich et al. (2018). This was then applied to obtain generalization bounds for the task of learning Markov-chains in Truong (2022a), however the generalization error was measured via the marginal cost and the $(\mu,c)$-RC type inequality was only applied for time-invariant neural networks. In contrast, we prove that along with time invariant models, stable SSMs, defined between certain Banach spaces, also satisfy equation (2) and apply it to deep structures.
In a recent work Trauger and Tewari (2024), the authors consider Transformers and implicitly establish similar inequalities to (2) by bounding different kinds of operator norms of the model and managed to extend it to a stack of Transformer layers. Besides these similarities, some key differences in our work are that Definition 10 provides an explicit way to combine SSMs with neural networks, even in residual blocks; we do not assume the SSM matrices to be bounded, instead we require the system norm to be bounded via stability, which is a weaker condition; and we upper bound the Rademacher complexity directly instead of bounding the covering number.
Interpretation of equation (2). Usually, the machine learning literature deals with functions between vector spaces, whereas in the above definition we consider functions between normed spaces. As a result, in order to apply this definition to any component of a deep SSM model, which are defined as maps between vector spaces in section 3.1, we need to equip the domain and range of these maps with some norms. The choice of these norms is arbitrary, but the constants $\mu$ and $c$ depend on it, and the analysis of a deep structure componentwise requires that the image of a middle layer is equiepped with the same norm as the domain of the succeeding layer. Foreshadowing the next section, this is exactly what we would like to do, namely show, that each component of a deep SSM model parametrized by some appropriate sets of parameters, is $(\mu,c)$-RC for some $\mu$ and $c$.
In case of affine layers and SSMs, this is proven in Lemma 13 and those results hold for unbounded $X_{1}$ and $X_{2}$ sets with $c=0$. Lemma 16 gives similar results for sigmoid and ReLU MLPs with $c>0$, which are also valid for ubounded input space. As for the GLU layers, it is possible to prove a $(\mu,c)$-RC property, but the constants depend on the size of $X_{1}$. Consequently, we restrict the input space of the nonlinear layers to a ball of a fixed radius. As these layers are preceded and succeeded by other components, we need to ensure that if the input of the deep SSM is from a ball of some radius $r$, the output of each layer is also in a ball of a radius that may depend on $r$ and the layer’s parameter set.
The following Lemma provides the way to establish a (2) type inequality for deep structures whose components are each $(\mu,c)$-RC.
Lemma 11 (Composition lemma)
Let $\Phi_{1}=\{\varphi_{1}:X_{1}\to X_{2}\}$ be $(\mu_{1},c_{1})$-RC and $\Phi_{2}=\{\varphi_{2}:X_{2}\to X_{3}\}$ be a $(\mu_{2},c_{2})$-RC. Then the set of compositions $\Phi_{2}\circ\Phi_{1}:=\{\varphi_{2}\circ\varphi_{1}:X_{1}\to X_{3}\mid\varphi_{1}\in\Phi_{1},\varphi_{2}\in\Phi_{2}\}$ is a $(\mu_{1}\mu_{2},\mu_{2}c_{1}+c_{2})$-RC.
The proof is in Appendix B. The upcoming corollary is straightforward by induction along with the fact that the pooling layer $f^{\text{Pool}}$ is $(1,0)$-RC (see Lemma 13).
Corollary 12
Let $f$ be a deep SSM model, i.e. according to Definition 6, $f=f^{\text{Dec}}\circ f^{\text{Pool}}\circ f^{\text{B}_{L}}\circ\ldots\circ f^{\text{B}_{1}}\circ f^{\text{Enc}}$, such that $f^{\text{Enc}}$, $f^{\text{Dec}}$ and each $f^{\text{B}_{i}}$ are $(\mu_{0},c_{0})$-RC, $(\mu_{L+1},c_{L+1})$-RC and $(\mu_{i},c_{i})$-RC respectively for all $i$. Then $\psi$ is $\left(\prod\limits_{i=0}^{L+1}\mu_{i},\sum\limits_{j=1}^{L}\left[\prod\limits_{i=j+1}^{L+1}\mu_{i}\right]c_{j}\right)$-RC.
What we have left are the following. First, in light of the previous corollary, we need to show that each component of a deep SSM model is $(\mu,c)$-RC for some $\mu$ and $c$ w.r.t. compatible normed spaces. Second, we need to show that the Rademacher complexity of a $(\mu,c)$-RC model set are bounded in terms of $\mu$ and $c$. We start with the first one.
Lemma 13
Let $\mathcal{W}_{\text{Enc}}$, $\mathcal{W}_{\text{Dec}}$, and $\mathcal{E}$ denote some sets of parameters of some fixed Encoder, Decoder and SSM layers, respectively. The corresponding function sets are
  • $\mathcal{F}_{\text{Enc}}=\{f^{\text{Enc}}=\langle W^{\text{Enc}},\cdot\rangle\mid W\in\mathcal{W}_{\text{Enc}}\}$,
  • $\mathcal{F}_{\text{Dec}}=\{f^{\text{Dec}}=\langle W^{\text{Dec}},\cdot\rangle\mid W\in\mathcal{W}_{\text{Dec}}\}$,
  • $\mathcal{F}_{\text{SSM}}=\{\mathcal{S}_{\Sigma}\mid\Sigma\in\mathcal{E}\}$,
where $\mathcal{S}_{\Sigma}$ denotes the input-output map of the dynamical system $\Sigma$. Then all of these function sets are $(\mu,c)$-RC according to the following table, where for any Banach space $\mathcal{X}$, $B_{\mathcal{X}}(t)=\{x\in\mathcal{X}\mid\left\lVert x\right\rVert_{\mathcal{X}}\leq r\}$ denotes the ball of radius $r$ centered in zero for arbitrary $r$.
$\mu$ $c$ $X_{1}$ $X_{2}$
$\mathcal{F}_{\text{Enc}}$ $\sup\limits_{W^{\text{Enc}}\in\mathcal{W}_{\text{Enc}}}\left\lVert W^{\text{Enc}}\right\rVert_{2,2}<K_{\text{Enc}}$ 0 $B_{\ell^{2,2}(\mathbb{R}^{n_{\text{in}}})}(r)$ $B_{\ell^{2,2}(\mathbb{R}^{n_{u}})}(K_{\text{Enc}}r)$
$\mathcal{F}_{\text{Dec}}$ $\sup\limits_{W^{\text{Dec}}\in\mathcal{W}_{\text{Dec}}}\left\lVert W^{\text{Dec}}\right\rVert_{\infty,\infty}<K_{\text{Dec}}$ 0 $B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r)$ $B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{\text{out}}})}(K_{\text{Dec}}r)$
$\sup\limits_{W^{\text{Dec}}\in\mathcal{W}_{\text{Dec}}}\left\lVert W^{\text{Dec}}\right\rVert_{\infty,\beta}<K_{\text{Dec}}$ 0 $B_{(\mathbb{R}^{n_{u}},\left\lVert\cdot\right\rVert_{\infty})}(r)$ $B_{(\mathbb{R}^{n_{\text{out}}},\left\lVert\cdot\right\rVert_{\beta})}(K_{\text{Dec}}r)$
$\mathcal{F}_{\text{SSM}}$ $\sup\limits_{\Sigma\in\mathcal{E}}\left\lVert\Sigma\right\rVert_{2}<K_{2}$ 0 $B_{\ell^{2,2}(\mathbb{R}^{n_{u}})}(r)$ $B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{y}})}(K_{2}r)$
$\sup\limits_{\Sigma\in\mathcal{E}}\left\lVert\Sigma\right\rVert_{1}<K_{1}$ 0 $B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r)$ $B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{y}})}(K_{1}r)$
Furthermore, the operation of $f^{\text{Pool}}$, defined in Definition 6, is $(1,0)$-RC between $X_{1}=B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r)$ and $X_{2}=B_{(\mathbb{R}^{n_{u}},\left\lVert\cdot\right\rVert_{\infty})}(r)$.
The proof is in Appendix B. We can see that the SSM layer can only increase the input’s complexity by the factor $\left\lVert\Sigma\right\rVert_{p}$, $p=1,2$, a quantity that gets smaller as the system gets more stable. This gets even more crucial when dealing with long range sequences, because the Neural Network layers are constant in time.
Remark 14
The results of of Lemma 13 hold for unbounded input spaces as well. The reason for restricting the input space to a ball of radius $r$ is the composition with MLPs or GLU layers, as discussed in the interpretation of Definition 10.
The last factor that influences the complexity are the Neural Network layers that follow the SSM layers in each SSM block (Definition 5). There are several results in the literature establishing upper bounds on the Rademacher complexity of MLPs and its variants. Additionally, in some cases the proof techniques directly imply the $(\mu,c)$-RC property of these models. As a result, we do not pay special attention to the MLP layers. We prove the $(\mu,c)$-RC property of GLU layers as they are commonly used in deep SSM structures, and deduct the values of $\mu$ and $c$ from previously known results in case of deep MLPs.
Lemma 15
Let $\mathcal{F}_{\text{GLU}}$ denote a set of GLU layers, i.e. $\mathcal{F}_{\text{GLU}}=\{GLU_{W}(\mathbf{u})=(GELU[\mathbf{u}])\odot\sigma(W(GELU(\mathbf{u})))\mid W\in\mathcal{W}\}$. Under the Bounded Input assumption in Assumption 9 we have that $\mathcal{F}_{\text{GLU}}$ is $(\mu,0)$-RC w.r.t. the spaces $X_{1}=B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}(K_{\mathbf{u}})=X_{2}$, where $\mu=16\left(K_{\mathbf{u}}\cdot\max\left\{\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{\infty,\infty},1\right\}+1\right)\left(\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{\infty,\infty}+1\right)$.
The proof is in Appendix B.
Lemma 16
Let $\mathcal{F}^{\rho}$ denote a set of hidden layers, i.e. =$\mathcal{F}=\{f_{W,b}(\mathbf{u})=\rho(W\mathbf{x}+\mathbf{b})\mid(W,\mathbf{b})\in\mathcal{W}\times\mathcal{B}\}$ and $\rho$ is either the sigmoid or the ReLU activation function. Let us assume that there exist $K_{W}>0$ and $K_{\mathbf{b}}>0$ such that $\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{\infty,\infty}<K_{W}$ and $\sup\limits_{\mathbf{b}\in\mathcal{B}}\left\lVert\mathbf{b}\right\rVert_{\infty}<K_{\mathbf{b}}$. Under the Bounded Input assumption in Assumption 9 we have that $\mathcal{F}^{ReLU}$ is $(\mu,c)$-RC w.r.t. the spaces $X_{1}=B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r)$, $X_{2}=B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{v}})}(K_{W}r+K_{\mathbf{b}})$, $\mu=4K_{W}$ and $c=4K_{W}K_{\mathbf{b}}$.
Under the extra assumption that the parameter set is symmetric to the origin, meaning that $(W,b)\in\mathcal{W}\times\mathcal{B}$ implies $(-W,-b)\in\mathcal{W}\times\mathcal{B}$, we have that $\mathcal{F}^{sigmoid}$ is $(\mu,c)$-RC w.r.t. the spaces $X_{1}=B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r)$, $X_{2}=B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{v}})}(1)$, $\mu=K_{W}$ and $c=K_{W}(K_{\mathbf{b}}+0.5)$.
The proof is in Appendix B. Combining the previous Lemma with Lemma 11 and Lemma 21 we get the following for deep networks.
Corollary 17
Let $g_{W,\mathbf{b}}(x)=W\mathbf{x}+\mathbf{b}$ represent an affine transformation and consider a set of deep networks $\mathcal{F}^{\rho}=\{f=f^{\rho}_{W_{1},\mathbf{b}_{1}}\circ\ldots\circ f^{\rho}_{W_{L},\mathbf{b}_{L}}\circ g_{W_{L+1},\mathbf{b}_{L+1}}\mid(W_{i},\mathbf{b}_{i})\in\mathcal{W}_{i}\times\mathcal{B}_{i},1\leq i\leq L+1\}$, where $f^{\rho}_{W_{i},\mathbf{b}_{i}}$ are hidden layers with the same activations $\rho$ equal to either sigmoid or ReLU. Let $K_{W}>0$ and $K_{\mathbf{b}}>0$ such that $\sup\limits_{W\in\mathcal{W}_{i}}\left\lVert W\right\rVert_{\infty,\infty}<K_{W}$ and $\sup\limits_{\mathbf{b}\in\mathcal{B}_{i}}\left\lVert\mathbf{b}\right\rVert_{\infty}<K_{\mathbf{b}}$ hold for all $11\leq i\leq L+$ 1. Under the assumptions of Lemma 16 we have that $\mathcal{F}^{ReLU}$ is $(\mu,c)$-RC w.r.t. with $\mu=4(L+1)K_{W}$ and $c=4(L+1)K_{W}K_{\mathbf{b}}\frac{L(L+1)}{2}$ and $\mathcal{F}^{sigmoid}$ is $(\mu,c)$-RC w.r.t. with $\mu=(L+1)K_{W}$ and $c=(L+1)K_{W}(K_{\mathbf{b}}+0.5)\frac{L(L+1)}{2}$.
So far we showed that each component of a deep SSM model satisfies Definition 10. We also proved that the composition of such components also satisfies the definition. The main theorem summarizes these results and exploits the fact that the Rademacher complexity of a $(\mu,c)$-RC set of models is upper bounded by terms containing $\mu$ and $c$.
Theorem 18
Let $\mathcal{F}$ be a set of deep SSM models, namely let $f\in\mathcal{F}$ has the form $f=f^{\text{Dec}}\circ f^{\text{Pool}}\circ f^{\text{B}_{L}}\circ\ldots\circ f^{\text{B}_{1}}\circ f^{\text{Enc}}$ with layer parameter sets $\mathcal{W}_{\text{Dec}}$, $\mathcal{W}_{\text{B}_{L}},\ldots,\mathcal{W}_{\text{B}_{1}}$ and $\mathcal{W}_{\text{Enc}}$ respectively, where $f^{\text{B}_{i}}$ is an SSM block for all $i$, i.e. $f^{\text{B}_{i}}(\mathbf{z})=g_{i}(\mathcal{S}_{\Sigma_{i}}(\mathbf{z})[k])+\alpha_{i}\mathbf{z}[k]$ and $\mathcal{W}_{\text{B}_{i}}=\mathcal{E}_{i}\times\mathcal{W}_{g_{i}}$. If Assumption 9 holds and each set of nonlinearities $g_{i}:X_{i}\rightarrow\hat{X}_{i}$ is $(\mu_{g_{i}},c_{g_{i}})$-RC, with $X_{1}=B_{\left(\mathbb{R}^{n_{u}},\left\lVert\cdot\right\rVert_{2}\right)}(K_{2}K_{\mathbf{u}}K_{\text{Enc}})$ $\hat{X}_{i}=B_{\left(\mathbb{R}^{n_{u}},\left\lVert\cdot\right\rVert_{\infty}\right)}(\hat{r}_{i})$ $X_{i+1}=B_{\left(\mathbb{R}^{n_{u}},\left\lVert\cdot\right\rVert_{\infty}\right)}(r_{i+1})$, $r_{i+1}\geq K_{1}(\hat{r}_{i}+|\alpha_{i}|r_{i})$, where $r_{1}=K_{\mathbf{u}}$, $i=1,\ldots,L$, then the following holds with probability at least $1-\delta$.
$\mathbb{P}_{S\sim\mathcal{D}^{N}}\left[\forall f\in\mathcal{F}\quad\mathcal{L}(f)-\mathcal{L}_{emp}^{S}(f)\leq\frac{\mu K_{\mathbf{u}}L_{l}+cL_{l}}{\sqrt{N}}+K_{l}\sqrt{\frac{2\log(4/\delta)}{N}}\right],$ (3)
$\mu\leq K_{\text{Enc}}K_{\text{Dec}}\left(\mu_{g_{1}}K_{2}+\alpha_{1}\right)\prod\limits_{i=2}^{L}\left(\mu_{g_{i}}K_{1}+\alpha_{i}\right)$, $c\leq K_{\text{Dec}}\sum\limits_{j=1}^{L}\left[\prod\limits_{i=j+1}^{L}\left(\mu_{g_{i}}K_{1}+\alpha_{i}\right)\right]c_{g_{j}}$ and $K_{l}>0$ such that $|l(\cdot,\cdot)|\leq K_{l}$. In particular, we obtain $K_{l}\leq 2L_{l}\max\{K_{\text{Dec}}r_{L+2},K_{y}\}$.
The proof can be found in Appendix B.
Appendix A Rademacher complexity
Definition 19
(Shalev-Shwartz and Ben-David, 2014, Def. 26.1) The Rademacher complexity of a bounded set $\mathcal{A}\subset\mathbb{R}^{m}$ is defined as
$R(\mathcal{A})=\mathbb{E}_{\mathbf{\sigma}}\Bigg{[}\sup_{a\in\mathcal{A}}\frac{1}{m}\sum\limits_{i=1}^{m}\sigma_{i}a_{i}\Bigg{]},$
where the random variables $\mathbf{\sigma}_{i}$ are i.i.d such that $\mathbb{P}[\sigma_{i}=1]=\mathbb{P}[\sigma=-1]=0.5$. The Rademacher complexity of a set of functions $\mathcal{F}$ over a set of samples $S=\{s_{1}\dots s_{m}\}$ is defined as $R_{S}(\mathcal{F})=R(\left\{(f(s_{1}),\dots,f(s_{m}))\mid f\in\mathcal{F}\right\}).$
The following is a standard theorem we use in the proof.
Theorem 20
(Shalev-Shwartz and Ben-David, 2014, Thm. 26.5) Let $L_{0}$ denote the set of functions of the form $(\mathbf{u},y)\mapsto l(f(\mathbf{u}),y)$ for $f\in\mathcal{F}$. Let $K_{l}$ be such that the functions from $L_{0}$ all take values from the interval $[0,K_{l}]$. Then for any $\delta\in(0,1)$ we have
$\mathbb{P}_{\mathbf{S}\sim\mathcal{D}^{N}}\Bigg{(}\forall f\in\mathcal{F}:\mathcal{L}(f)-\mathcal{L}^{S}_{emp}(f)\leq 2R_{S}(L_{0})+K_{l}\sqrt{\frac{2\log(4/\delta)}{N}}\Bigg{)}\geq 1-\delta.$
Appendix B Proofs
In this section we need to prove $(\mu,c)$-RC property for linear (or affine) transformations which are constant in time, in many cases. For better readibility, we only do the calculations once and use it as a lemma.
Lemma 21
Let $\mathbf{u}\in\ell^{p,p}(\mathbb{R}^{n_{u}})=:X_{1}$ and let $f(\mathbf{u})=f_{W,\mathbf{b}}(\mathbf{u}):=W\mathbf{u}+\mathbf{b}\in\ell^{q,q}(\mathbb{R}^{n_{v}})=:X_{2}$, where $W\in\mathbb{R}^{n_{v}\times n_{u}},\mathbf{b}\in\mathbb{R}^{n_{v}}$, and by definition $W\mathbf{u}\in X_{2}$ such that $(W\mathbf{u}+b)[k]=W\mathbf{u}[k]+\mathbf{b}$, i.e. it is an affine transformation constant in time. We consider the cases
  • a)
    $p=q=2$,
  • b)
    $p=2,q=\infty$,
  • c)
    $p=q=\infty$.
Let us assume that there exist constants $K_{W}$ and $K_{\mathbf{b}}$ such that $\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{p,q}\leq K_{W}$ and $\sup\limits_{\mathbf{b}\in\mathcal{B}}\left\lVert\mathbf{b}\right\rVert_{q}\leq K_{\mathbf{b}}$. Then the set of transformation =$\mathcal{F}=\{f_{W,\mathbf{b}}\mid W\in\mathcal{W},\mathbf{b}\in\mathcal{B}\}$ is $(K_{W},K_{\mathbf{b}})$-RC w.r.t. $X_{1}$ and $X_{2}$. Furthermore, the image of the ball $B_{X_{1}}(r)$ under $f\in\mathcal{F}$ is contained in $B_{X_{2}}(K_{W}r+K_{\mathbf{b}})$, and therefore $\mathcal{F}|_{B_{X_{1}}(r)}$ is also $(K_{W},K_{\mathbf{b}})$-RC w.r.t the spaces $B_{X_{1}}(r)$ and $B_{X_{2}}(K_{W}r+K_{\mathbf{b}})$.
Proof  First let us prove a simple fact about Rademacher random variables that we will need, namely if $\sigma=\{\sigma_{i}\}_{i=1}^{N}$ is a sequence of i.i.d. Rademacher variables, then
$\mathbb{E}_{\sigma}\left[\left|\sum\limits_{i=1}^{N}\sigma_{i}\right|\right]\leq\sqrt{N}.$ (4)
This is true, because
$\mathbb{E}_{\sigma}\left[\left|\sum\limits_{i=1}^{N}\sigma_{i}\right|\right]=\sqrt{\left(\mathbb{E}_{\sigma}\left[\left|\sum\limits_{i=1}^{N}\sigma_{i}\right|\right]\right)^{2}}\leq\sqrt{\mathbb{E}_{\sigma}\left[\left|\sum\limits_{i=1}^{N}\sigma_{i}\right|^{2}\right]}$
$=\sqrt{\mathbb{E}_{\sigma}\left[\sum\limits_{i=1}^{N}\sigma^{2}_{i}+2\sum\limits_{i,j=1}^{N}\sigma_{i}\sigma_{j}\right]}=\sqrt{\sum\limits_{i=1}^{N}\mathbb{E}_{\sigma}\left[\sigma^{2}_{i}\right]+2\sum\limits_{i,j=1}^{N}\mathbb{E}_{\sigma}\left[\sigma_{i}\sigma_{j}\right]}=\sqrt{N},$
where the first inequality follows from Jensen’s inequality and the last equality follows from the linearity of the expectation, and the facts that $\sigma_{i}$ are Rademacher variables and form and i.i.d sample.
Let $Y\in\{X_{1},B_{X_{1}}(r)\}$.
a) Let $\underline{\mathbf{b}}\in Y$ denote the sequence for which $\underline{\mathbf{b}}[k]=\mathbf{b}$ for all $k\in[T]$. For any $Z\subseteq Y$ we have
$\mathbb{E}_{\sigma}\left[\sup\limits_{(W,\mathbf{b})\in\mathcal{W}\times\mathcal{B}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}(W\mathbf{u}_{i}+\mathbf{b})\right\rVert_{Y}\right]$
$\leq\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}W\mathbf{u}_{i}\right\rVert_{Y}\right]+\mathbb{E}_{\sigma}\left[\sup\limits_{\mathbf{b}\in\mathcal{B}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\underline{\mathbf{b}}\right\rVert_{Y}\right]$
$=\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\sqrt{\sum\limits_{k=1}^{T}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}W\mathbf{u}_{i}[k]\right\rVert_{2}^{2}}\right]+\mathbb{E}_{\sigma}\left[\left|\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\right|\sup\limits_{\mathbf{b}\in\mathcal{B}}\left\lVert\underline{\mathbf{b}}\right\rVert_{Y}\right]$
$\leq\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\sqrt{\sum\limits_{k=1}^{T}\left\lVert W\left(\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}[k]\right)\right\rVert_{2}^{2}}\right]+\frac{1}{\sqrt{N}}\sup\limits_{\mathbf{b}\in\mathcal{B}}\left\lVert\mathbf{b}\right\rVert_{2}$
$\leq\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert W\right\rVert_{2,2}\sqrt{\sum\limits_{k=1}^{T}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}[k]\right\rVert_{2}^{2}}\right]+\frac{1}{\sqrt{N}}\sup\limits_{\mathbf{b}\in\mathcal{B}}\left\lVert\mathbf{b}\right\rVert_{2}$
$=\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{2,2}\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\right\rVert_{Y}\right]+\frac{1}{\sqrt{N}}\sup\limits_{\mathbf{b}\in\mathcal{B}}\left\lVert\mathbf{b}\right\rVert_{2},$
where the first inequality follows from the triangle inequality, the first equality is the definition of the norm, the second inequality follows from the equation (4) and the linearity of $W$, and the last inequality is a standard inequality for matrix norms.
b) The first inequalities referring to the bias term are exactly the same and hold for the infinity norm as well as in case a). We only have to deal with the term
$\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}W\mathbf{u}_{i}\right\rVert_{Y}\right],$
but now $\mathbf{u}_{i}\in\ell^{2,2}(\mathbb{R}^{n_{u}})$ and $W\mathbf{u}_{i}\in\ell^{\infty,\infty}(\mathbb{R}^{n_{v}})$. We have
$\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}W\mathbf{u}_{i}\right\rVert_{Y}\right]=\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert W\left(\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\right)\right\rVert_{Y}\right]$
$=\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\sup\limits_{1\leq k\leq T}\left\lVert W\left(\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}[k]\right)\right\rVert_{\infty}\right]$
$\leq\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\sup\limits_{1\leq k\leq T}\left\lVert W\right\rVert_{2,\infty}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}[k]\right\rVert_{2}\right]$
$=\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{2,\infty}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\sqrt{\sup\limits_{1\leq k\leq T}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}[k]\right\rVert_{2}^{2}}\right]$
$\leq\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{2,\infty}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\sqrt{\sum\limits_{k=1}^{T}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}[k]\right\rVert_{2}^{2}}\right]$
$\leq\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{2,\infty}\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\right\rVert_{\ell^{2,2}(\mathbb{R}^{n_{u}})}\right],$
where the second equality is the definition of the norm, the first inequality is the standard inequality for induced matrix norms.
c) Again, we only need to deal with the term
$\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}W\mathbf{u}_{i}\right\rVert_{Y}\right],$
but now $\mathbf{u}_{i}\in\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})$ and $W\mathbf{u}_{i}\in\ell^{\infty,\infty}(\mathbb{R}^{n_{v}})$. We have
$\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}W\mathbf{u}_{i}\right\rVert_{Y}\right]=\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert W\left(\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\right)\right\rVert_{Y}\right]$
$=\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\sup\limits_{1\leq k\leq T}\left\lVert W\left(\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}[k]\right)\right\rVert_{\infty}\right]$
$\leq\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\sup\limits_{1\leq k\leq T}\left\lVert W\right\rVert_{\infty,\infty}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}[k]\right\rVert_{\infty}\right]$
$=\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{\infty,\infty}\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right],$
where the inequality is again the standard inequality for induced matrix norms.
We can see that the calculactions hold if the transformations are restricted to the ball $B_{X_{1}}(r)$ for any choice of $X_{1}$ we consider. The radius can grow as
$\left\lVert W\mathbf{u}+\mathbf{b}\right\rVert_{\ell^{q,q}(\mathbb{R}^{n_{v}})}\leq\left\lVert W\mathbf{u}\right\rVert_{\ell^{q,q}(\mathbb{R}^{n_{v}})}+\left\lVert\underline{\mathbf{b}}\right\rVert_{\ell^{q,q}(\mathbb{R}^{n_{v}})}\leq\left\lVert W\right\rVert_{p,q}\left\lVert\mathbf{u}\right\rVert_{\ell^{p,p}(\mathbb{R}^{n_{u}})}+\left\lVert\mathbf{b}\right\rVert_{q}.$
Proof [Proof of Lemma 11] Let $Z\subseteq X_{1}^{N}$ and $\tilde{Z}=\{\{\varphi_{1}(\mathbf{u}_{i})\}_{i=1}^{N}\mid\varphi_{1}\in\Phi_{1}\}$. We have
$\mathbb{E}_{\sigma}\left[\sup\limits_{\varphi_{2}\in\Phi_{2}}\sup\limits_{\varphi_{1}\in\Phi_{1}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\varphi_{2}(\varphi_{1}(\mathbf{u}_{i}))\right\rVert_{X_{3}}\right]$
$=\mathbb{E}_{\sigma}\left[\sup\limits_{\varphi_{2}\in\Phi_{2}}\sup\limits_{\{\mathbf{v}_{i}\}_{i=1}^{N}\in\tilde{Z}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\varphi_{2}(\mathbf{v}_{i})\right\rVert_{X_{3}}\right]$
$\leq\mu_{2}\mathbb{E}_{\sigma}\left[\sup\limits_{\varphi_{1}\in\Phi_{1}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\varphi_{1}(\mathbf{u}_{i})\right\rVert_{X_{2}}\right]+\frac{c_{2}}{\sqrt{N}}$
$\leq\mu_{2}\mu_{1}\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\right\rVert_{X_{1}}\right]+\mu_{2}\frac{c_{1}}{\sqrt{N}}+\frac{c_{2}}{\sqrt{N}}$
Proof [Proof of Lemma 13]
Encoder and Decoder. The Encoder is case a), while the Decoder is case b) in Lemma 21.
SSM. As discussed in Section 3.1, an SSM is equivalent to a linear transformation called its input-output map. Therefore, by Lemma 21, the SSM is $(\mu,0)$-RC in both cases, where $\mu$ is the operator norm of the input-output map. Combining this with Lemma 3 yields the result.
Remark 22
As the value of $T$ is fixed, the input-output map can be described by the so-called Toeplitz matrix of the system. In this case, the operator norm equals to the appropriate induced matrix norm of the Toeplitz matrix. For the case of $T=\infty$, the input-output map still exists and is a linear operator. The proof of Lemma 21 holds in this case as well for operator norms.
Pooling. For any $Z\subseteq\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})$ we have
$\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}f^{\text{Pool}}(\mathbf{z}_{i})\right\rVert_{\infty}\right]$
$=\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\sup\limits_{1\leq j\leq n_{u}}\left|\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\left(\frac{1}{T}\sum\limits_{k=1}^{T}\mathbf{z}_{i}^{(j)}[k]\right)\right|\right]$
$=\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\sup\limits_{1\leq j\leq n_{u}}\left|\frac{1}{T}\sum\limits_{k=1}^{T}\left(\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}^{(j)}[k]\right)\right|\right]$
$\leq\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\frac{1}{T}\sum\limits_{k=1}^{T}\sup\limits_{1\leq j\leq n_{u}}\left|\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}^{(j)}[k]\right|\right]$
$=\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\frac{1}{T}\sum\limits_{k=1}^{T}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}[k]\right\rVert_{\infty}\right]$
$\leq\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
Proof [Proof of Lemma 15] First of all, we show that the function $h:(\mathbb{R}^{2},\left\lVert\cdot\right\rVert_{2})\to(\mathbb{R},|\cdot|)$ defined as $h(\mathbf{x})=x_{1}\cdot\sigma(x_{2})$ is $\sqrt{2}(K+1)$-Lipschitz on a bounded domain, where $|x_{i}|\leq K$ for all $\mathbf{x}\in\mathbb{R}^{2}$ we consider. We will later specify the value of $K$ in relation to Assumption 9. By the sigmoid being 1-Lipschitz, we have
$|h(\mathbf{x})-h(\mathbf{y})|=|x_{1}\sigma(x_{2})-y_{1}\sigma(x_{2})+y_{1}\sigma(x_{2})-y_{1}\sigma(y_{2})|\leq$
$|(x_{1}-y_{1})\sigma(x_{2})|+|y_{1}(\sigma(x_{2})-\sigma(y_{2}))|\leq|x_{1}-y_{1}|+|y_{1}||x_{2}-y_{2}|$
$\leq\sqrt{2}(K+1)\left\lVert\mathbf{x}-\mathbf{y}\right\rVert_{2}$
Second, we recall Corollary 4 in Maurer (2016).
Theorem 23 (Maurer (2016))
Let $\mathcal{X}$ be any set, $(\mathbf{x}_{1},\ldots,\mathbf{x}_{N})\in\mathcal{X}^{N}$, let $\mathcal{F}$ be a set of functions $f:\mathcal{X}\to\ell^{2}(\mathbb{R}^{m})$ and let $h:\ell^{2}(\mathbb{R}^{m})\to\mathbb{R}$ be an $L$-Lipschitz function. Under $f_{k}$ denoting the $k$-th component function of $f$ and $\sigma_{ik}$ being a doubly indexed Rademacher variable, we have
$\mathbb{E}_{\sigma}\left[\sup\limits_{f\in\mathcal{F}}\sum\limits_{i=1}^{N}\sigma_{i}h(f(\mathbf{x}_{i}))\leq\sqrt{2}L\mathbb{E}_{\sigma}\left[\sup\limits_{f\in\mathcal{F}}\sum\limits_{i=1}^{N}\sum\limits_{k=1}^{m}\sigma_{ik}f_{k}(\mathbf{x}_{i})\right]\right].$
We wish to apply Theorem 23 to GLU layers. For any $Z\subseteq\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})$ we have
$\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}GLU_{W}(\mathbf{z}_{i})\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
$=\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\sup\limits_{1\leq k\leq T}\sup\limits_{1\leq j\leq n_{u}}\left|\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}GLU_{W}^{(j)}(\mathbf{z}_{i})[k]\right|\right]$
Now this is an alternative version of the Rademacher complexity, where we take the absolute value of the Rademacher average. In order to apply Theorem 23, we reduce the problem to the usual Rademacher complexity. In turn, we can apply the last chain of inequalities in the proof of Proposition 6.2 in Hajek and Raginsky (2019). Concretely, by denoting $\mathbf{O}=\{\mathbf{0}\}_{i=1}^{N}$ and noticing that $GLU_{W}(0)=0$, we have
$\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\sup\limits_{1\leq k\leq T}\sup\limits_{1\leq j\leq n_{u}}\left|\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}GLU_{W}^{(j)}(\mathbf{z}_{i})[k]\right|\right]$
$\leq 2\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z\cup\{\mathbf{O}\}}\sup\limits_{1\leq k\leq T}\sup\limits_{1\leq j\leq n_{u}}\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}GLU_{W}^{(j)}(\mathbf{z}_{i})[k]\right]$
Let $\mathbf{x}_{i}=i$, $i=1,\ldots,N$ and let =$\mathcal{F}=\{f_{W,\underline{z},k,j}\mid(W,\underline{z},k,j)\in\mathcal{W}\times(Z\cup\{0\})\times[T]\times[n_{u}]\}$ such that $f_{W,\underline{z},k,j}(\mathbf{x}_{i})=\begin{bmatrix}GELU(\mathbf{z}_{i}[k])^{(j)}&(W(GELU(\mathbf{z}_{i}[k])))^{(j)}\end{bmatrix}^{T}$ for $\underline{z}=\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z$. Since $Z\subseteq(B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}(K_{\mathbf{u}}))^{N}$, it follows for all $\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z$ and for all $k\in\mathbb{N}$ that $\|\mathbf{z}_{i}[k]\|_{\infty}\leq K_{\mathbf{u}}$, and hence $|GELU(\mathbf{z}_{i}[k])^{(j)}|<K_{\mathbf{u}}$, leading to $|W(GELU(\mathbf{z}_{i}[k]))^{(j)}|<\sup\limits_{W\in\mathcal{W}}\|W\|_{\infty,\infty}\cdot K_{\mathbf{u}}$. In particular, $GLU^{(j)}_{W}(\mathbf{z}_{i})[k]=h(f_{W,\underline{z},k,j}(\mathbf{x}_{i}))=h|_{B}(f_{W,\underline{z},k,j}(\mathbf{x}_{i}))$, where $h|_{B}$ is the restriction of $h$ to $B=\{x\in\mathbb{R}^{2}\mid\|x\|_{\infty}<K\}$, and hence $h|_{B}$ is $\sqrt{2}(K+1)$-Lipschitz. Therefore we can set $K=\max\{K_{\mathbf{u}},\sup\limits_{W\in\mathcal{W}}\|W\|_{\infty,\infty}\cdot K_{\mathbf{u}}\}$.
We are ready to apply Theorem 23, together with the GLU definition and its $\sqrt{2}(K+1)$-Lipschitzness, we have
$2\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z\cup\{\mathbf{O}\}}\sup\limits_{1\leq k\leq T}\sup\limits_{1\leq j\leq n_{u}}\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}GLU_{W}^{(j)}(\mathbf{z}_{i})[k]\right]$
$=2\mathbb{E}_{\sigma}\left[\sup\limits_{f\in\mathcal{F}}\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}f(\mathbf{x}_{i})\right]\leq 4(K+1)\underbrace{\mathbb{E}_{\sigma}\left[\sup\limits_{f\in\mathcal{F}}\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}GELU(\mathbf{z}_{i}[k])^{(j)}\right]}_{A}$
$+4(K+1)\underbrace{\mathbb{E}_{\sigma}\left[\sup\limits_{f\in\mathcal{F}}\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}W(GELU(\mathbf{z}_{i}))^{(j)}[k]\right]}_{B}$
Due to the definition of GELU, its 2-Lipschitzness Qi et al. (2023) and (Ledoux and Talagrand, 1991, Theorem 4.12) we have
$A=\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z\cup\{\mathbf{O}\}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}GELU(\mathbf{z}_{i})\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]=$
$\leq 4\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z\cup\{\mathbf{O}\}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]=4\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
and
$B=\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in\{\mathbf{O}\}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}W(GELU(\mathbf{z}_{i}))\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
$\leq 4\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{\infty,\infty}\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z\{\mathbf{O}\}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
$=4\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{\infty,\infty}\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
Here we used the linearity of $W$ and the exact same calculation as in the proof of Lemma 21.
By combining the inequalities above, it follows that
$\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\sup\limits_{1\leq k\leq T}\sup\limits_{1\leq j\leq n_{u}}\left|\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}GLU_{W}^{(j)}(\mathbf{z}_{i})[k]\right|\right]\leq$
$16(K+1)\left(\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{\infty,\infty}+1\right)\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
Substituting the value of $K$ gives the result.
Proof [Proof of Lemma 16]
MLP with sigmoid activations. Consider a single hidden layer $f(\mathbf{x})=\rho(g(\mathbf{x}))$, where $g(x)=W\mathbf{x}+\mathbf{b}$ is the preactivation and let $\mathcal{G}=\left\{g(x)=W\mathbf{x}+\mathbf{b}\mid W\in\mathcal{W},\mathbf{b}\in\mathcal{B}\right\}$ denote the set of possible preactivation functions. As compared to Definition 7, we omit the subscript from the notation of $g$. For an input sequence $\mathbf{z}\in\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})$ let $g(\mathbf{z})\in\ell^{\infty,\infty}(\mathbb{R}^{n_{v}})$ mean that we apply $g$ for each timestep independently, i.e. $g(\mathbf{z})[k]=g(\mathbf{z}[k])$. We have
$\mathbb{E}_{\sigma}\left[\sup\limits_{g\in\mathcal{G}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\rho(g(\mathbf{z}_{i}))\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
$=\mathbb{E}_{\sigma}\left[\sup\limits_{(W,\mathbf{b})\in\mathcal{W}\times\mathcal{B}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\sup\limits_{1\leq k\leq T}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\rho(W\mathbf{z}_{i}[k]+\mathbf{b})\right\rVert_{\infty}\right]$
Let $\mathbf{x}_{i}=i$, $i=1,\ldots,N$ and let =$\mathcal{H}=\{h_{W,\mathbf{b},\underline{z},k}\mid(W,\mathbf{b},\underline{z},k)\in\mathcal{W}\times\mathcal{B}\times(Z\cup\{0\})\times[T]\}$ such that $h_{W,\mathbf{b},\underline{z},k}(\mathbf{x}_{i})=g(\mathbf{z}_{i}[k])$. Under the assumption that $\mathcal{H}$ is symmetric to the origin, meaning that $h\in\mathcal{H}$ implies $-h\in\mathcal{H}$ (equivalently $(W,\mathbf{b})\in\mathcal{W}\times\mathcal{B}$ implies $(-W,-\mathbf{b})\in\mathcal{W}\times\mathcal{B}$), we can apply (Truong, 2022b, Theorem 9) for the sigmoid activation and hence $\rho-\rho(0)$ being odd, as follows.
$\mathbb{E}_{\sigma}\left[\sup\limits_{(W,\mathbf{b})\in\mathcal{W}\times\mathcal{B}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\sup\limits_{1\leq k\leq T}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\rho(W\mathbf{z}_{i}[k]+\mathbf{b})\right\rVert_{\infty}\right]$
$=\mathbb{E}_{\sigma}\left[\sup\limits_{h\in\mathcal{H}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\rho(h(\mathbf{x}_{i}))\right\rVert_{\infty}\right]$
$\leq\mathbb{E}_{\sigma}\left[\sup\limits_{h\in\mathcal{H}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}h(\mathbf{x}_{i})\right\rVert_{\infty}\right]+\frac{1}{2\sqrt{N}}$
$=\mathbb{E}_{\sigma}\left[\sup\limits_{(W,\mathbf{b})\in\mathcal{W}\times\mathcal{B}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\sup\limits_{1\leq k\leq T}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}(W\mathbf{z}_{i}[k]+\mathbf{b})\right\rVert_{\infty}\right]+\frac{1}{2\sqrt{N}}$
$=\mathbb{E}_{\sigma}\left[\sup\limits_{(W,b)\in\mathcal{W}\times\mathcal{B}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}(W\mathbf{z}_{i}+\mathbf{b})\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{v}})}\right]+\frac{1}{2\sqrt{N}},$
because the sigmoid is 1-Lipschitz and $\rho(0)=0.5$. Now we can apply Lemma 21 to get that
$\mathbb{E}_{\sigma}\left[\sup\limits_{(W,\mathbf{b})\in\mathcal{W}\times\mathcal{B}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}(W\mathbf{z}_{i}+\mathbf{b})\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{v}})}\right]+\frac{1}{2\sqrt{N}}$
$\leq\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{\infty,\infty}\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]+\frac{1}{\sqrt{N}}\sup\limits_{\mathbf{b}\in\mathcal{B}}\left\lVert\mathbf{b}\right\rVert_{\infty}+\frac{1}{2\sqrt{N}}$
Therefore, the sigmoid MLP layer is $(K_{W},K_{\mathbf{b}}+0.5)$-RC. The restriction of an MLP to the ball $B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r)$ maps to the ball $B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{v}})}(1)$, because of the elementwise sigmoid activation.
MLP with ReLU activations. Similarly to the sigmoid case, we assume the upper bounds $K_{W}$ and $K_{\mathbf{b}}$ exist, but we don’t assume the symmetry of the parameter set. The proof is the same as in the sigmoid case up to the first inequality. Here we can apply (Ledoux and Talagrand, 1991, Equation 4.20) (this is the same idea as in the proof of (Golowich et al., 2018, Lemma 2)) to get
$\mathbb{E}_{\sigma}\left[\sup\limits_{h\in\mathcal{H}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\rho(h(\mathbf{x}_{i}))\right\rVert_{\infty}\right]\leq 4\mathbb{E}_{\sigma}\left[\sup\limits_{h\in\mathcal{H}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}h(\mathbf{x}_{i})\right\rVert_{\infty}\right],$
where we used that $\rho(x)=ReLU(x)$ is 1-Lipschitz and the same logic for the alternative definition of the Rademacher complexity as in the proof of Lemma 15, which results in a constant factor of 2. The constant 4 is then obtained by the additional constant factor 2 from Talagrand’s lemma. The rest of proof is identical to the sigmoid case.
The restriction of an MLP to the ball $B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r)$ maps to the ball $B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{v}})}(K_{W}r+K_{\mathbf{b}})$, because the elementwise ReLU does not increase the infinity norm, hence we can apply Lemma 21.
Proof [Proof of Theorem 18] We wish to apply Theorem 20 to the set of deep SSM models $\mathcal{F}$. Let us fix a random sample $S=\{\mathbf{u}_{1},\ldots,\mathbf{u}_{N}\}\subset\left(\ell^{2,2}(\mathbb{R}^{n_{\text{in}}})\right)^{N}$. As the loss function is Lipschitz according to Assumption 9, we have that for any $f\in\mathcal{F}$
$|l(f(\mathbf{u}),y)|\leq 2L_{l}\max\{f(\mathbf{u}),y\}\leq 2L_{l}\max\{K_{\text{Dec}}r_{L+2},K_{y}\},$
thus $K_{l}\leq 2L_{l}\max\{K_{\text{Dec}}r_{L+2},K_{y}\}$.
Again by the Lipschitzness of the loss and the Contraction lemma (Shalev-Shwartz and Ben-David, 2014, Lemma 26.9) we have
$R_{S}(L_{0})\leq L_{l}\cdot R_{S}(\mathcal{F}).$
It is enough to bound the Rademacher complexity of $\mathcal{F}$ to conclude the proof. Let us consider the deep SSM models as a composite of mappings as
[…]
Therefore, the SSM layer in the first SSM block is considered as a map $B_{\ell^{2,2}(\mathbb{R}^{n_{u}})}(K_{\text{Enc}}K_{\mathbf{u}})\to B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r_{1})$, while the rest of the SSM layers in the SSM blocks are considered as a map $B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r_{i})\to B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r_{i+1})$.This is needed, because the Encoder is constant in time, therefore the Composition Lemma wouldn’t be able to carry the $\ell^{2,2}$ norm of the input through the chain of estimation along the entire model. This is one of the key technical points which makes it possible to establish a time independent bound.
By the conditions of the Theorem and the stability assumption in Assumption 9 we have that the Encoder, Decoder, Pooling, SSM and MLP layers are each $(\mu,c)$-RC for some $\mu$ and $c$ from Lemma 13. By Lemma 11 we have that the composition of an SSM layer and an MLP is $(\mu,c)$-RC. A residual SSM block is then $(\mu+\alpha,c)$-RC, because
$\mathbb{E}_{\sigma}\left[\sup\limits_{g\circ\mathcal{S}_{\Sigma}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}(g(\mathcal{S}_{\Sigma}(\mathbf{z}_{i}))+\alpha\mathbf{z}_{i})\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]\leq$
$\mathbb{E}_{\sigma}\left[\sup\limits_{g\circ\mathcal{S}_{\Sigma}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}g(\mathcal{S}_{\Sigma}(\mathbf{z}_{i}))\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]+\alpha\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
$\leq(\mu+\alpha)\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]+\frac{c}{\sqrt{N}}$
Hence, by Corollary 12, the whole deep SSM model is $(\mu,c)$-RC as a map between $X_{1}=B_{\ell^{2,2}(\mathbb{R}^{n_{\text{in}}})}(K_{\mathbf{u}})$ and $X_{2}=(\mathbb{R},|\cdot|)$. The Theorem is then a direct corollary of the following Lemma.
Lemma 24
Let $\mathcal{F}$ be a set of functions that is $(\mu,c)$-RC w.r.t. $X_{1}=B_{\ell^{2,2}(\mathbb{R}^{n_{\text{in}}})}(K_{\mathbf{u}})$ and $X_{2}=(\mathbb{R},|\cdot|)$. The Rademacher complexity of $\mathcal{F}$ w.r.t. some dataset $S$ for which Assumption 9 holds, admits the following inequality.
$R_{S}(\mathcal{F})\leq\frac{\mu K_{\mathbf{u}}+c}{\sqrt{N}}.$
Proof
$R_{S}(\mathcal{F})=R(\left\{(f(\mathbf{u}_{1}),\dots,f(\mathbf{u}_{N}))^{T}\mid f\in\mathcal{F}\right\})=\mathbb{E}_{\sigma}\left[\sup\limits_{f\in\mathcal{F}}\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}f(\mathbf{u}_{i})\right]$
$\leq\mu\mathbb{E}_{\sigma}\left[\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\right\rVert_{\ell^{2,2}(\mathbb{R}^{n_{\text{in}}})}\right]+\frac{c}{\sqrt{N}}$
By definition
$\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\right\rVert_{\ell^{2,2}(\mathbb{R}^{n_{\text{in}}})}=\sqrt{\sum\limits_{k=1}^{T}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}[k]\right\rVert^{2}_{2}}$
$=\sqrt{\sum\limits_{k=1}^{T}\left\langle\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}[k],\frac{1}{N}\sum\limits_{j=1}^{N}\sigma_{j}\mathbf{u}_{j}[k]\right\rangle_{\mathbb{R}^{n_{\text{in}}}}}$
$=\sqrt{\sum\limits_{k=1}^{T}\frac{1}{N^{2}}\sum\limits_{i=1}^{N}\sum\limits_{j=1}^{N}\sigma_{i}\sigma_{j}\left\langle\mathbf{u}_{i}[k],\mathbf{u}_{j}[k]\right\rangle_{\mathbb{R}^{n_{\text{in}}}}}$
Therefore
$\mathbb{E}_{\sigma}\left[\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\right\rVert_{\ell^{2,2}(\mathbb{R}^{n_{\text{in}}})}\right]=\mathbb{E}_{\sigma}\left[\sqrt{\sum\limits_{k=1}^{T}\frac{1}{N^{2}}\sum\limits_{i=1}^{N}\sum\limits_{j=1}^{N}\sigma_{i}\sigma_{j}\left\langle\mathbf{u}_{i}[k],\mathbf{u}_{j}[k]\right\rangle_{\mathbb{R}^{n_{\text{in}}}}}\right]$
$\leq\sqrt{\mathbb{E}_{\sigma}\left[\sum\limits_{k=1}^{T}\frac{1}{N^{2}}\sum\limits_{i=1}^{N}\sum\limits_{j=1}^{N}\sigma_{i}\sigma_{j}\left\langle\mathbf{u}_{i}[k],\mathbf{u}_{j}[k]\right\rangle_{\mathbb{R}^{n_{\text{in}}}}\right]}$
$=\sqrt{\sum\limits_{k=1}^{T}\frac{1}{N^{2}}\sum\limits_{i=1}^{N}\sum\limits_{j=1}^{N}\mathbb{E}_{\sigma}\left[\sigma_{i}\sigma_{j}\right]\left\langle\mathbf{u}_{i}[k],\mathbf{u}_{j}[k]\right\rangle_{\mathbb{R}^{n_{\text{in}}}}}$
$=\sqrt{\sum\limits_{k=1}^{T}\frac{1}{N^{2}}\sum\limits_{i=1}^{N}\mathbb{E}_{\sigma}\left[\sigma_{i}^{2}\right]\left\langle\mathbf{u}_{i}[k],\mathbf{u}_{i}[k]\right\rangle_{\mathbb{R}^{n_{\text{in}}}}}$
$=\sqrt{\frac{1}{N^{2}}\sum\limits_{i=1}^{N}\sum\limits_{k=1}^{T}\left\lVert\mathbf{u}_{i}[k]\right\rVert_{2}^{2}}=\sqrt{\frac{1}{N^{2}}\sum\limits_{i=1}^{N}\left\lVert\mathbf{u}_{i}\right\rVert^{2}_{\ell^{2,2}(\mathbb{R}^{n_{\text{in}}})}}\leq\sqrt{\frac{1}{N^{2}}NK_{\mathbf{u}}^{2}}\leq\frac{K_{\mathbf{u}}}{\sqrt{N}}$
Hence we have
$R_{S}(\mathcal{F})\leq\frac{\mu K_{\mathbf{u}}+c}{\sqrt{N}}$
The constants $\mu$ and $c$ are obtained by substituting the results of Lemma 13 into the Corollary 12.